OpenStudy (anonymous):

I need help with 5 problems please ! Its solving quadratic equations with complex solutions. Heres the first one: 1. Solve -3x2 - 4x - 4 = 0.

OpenStudy (anonymous):

they are all multiple choice..

OpenStudy (anonymous):

First multiple choice options: A.\[2\pm4i \sqrt{2}\over 3\] B.\[2\pm2i \sqrt{2} \over 3\] C.\[-2\pm 2i \sqrt{2} \over 3\] D.\[-2\pm 4i \sqrt{2} \over 3\]

OpenStudy (anonymous):

Just a quick question but is it 3x or 3*2

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

not 3 times 2 sorry for the confusion

OpenStudy (anonymous):



OpenStudy (anonymous):

this is one of four problems i need help with

OpenStudy (anonymous):

So how do you get the problem to go from that to the multiple choice answers above..

OpenStudy (anonymous):

It'll be A but it's really hard to convert

OpenStudy (anonymous):

/: okay .. can you help me with the other ones too?

OpenStudy (anonymous):

I sure can try

OpenStudy (anonymous):

thats all i can ask! Solve -3x2 + 30x - 90 = 0.

OpenStudy (anonymous):

Did you see how I did the first one just do that post it then I'll help u convert it

OpenStudy (anonymous):

Options: A.\[5\pm2i\] B.\[5\pm i \sqrt{5}\] C.\[10\pm 2i \sqrt{5}\] D.\[10\pm i \sqrt{5}\]

OpenStudy (anonymous):

how did you get the 12 did you multiply the 2 and 4 or howd you get that

OpenStudy (anonymous):

See you have two sets of numbers with a variable so you multiply them, or add them

OpenStudy (anonymous):

okay so for the one i just posted..id multiply the -3x^2 and 30x?

OpenStudy (anonymous):

Wait its add but the answer to the first one is still A

OpenStudy (anonymous):

okay ! so for the second one we add the 3x^2 and 30x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so its 33x

OpenStudy (anonymous):

okay and then after you add what do you do

OpenStudy (anonymous):

its actually 27 because its a -3 is that correct?

OpenStudy (anonymous):

@sean_davis0821

OpenStudy (anonymous):

Oh yes u r right

OpenStudy (anonymous):

and then after this step?

OpenStudy (anonymous):

Sorry I about messed u up

OpenStudy (anonymous):

its no problem. a simple error